OpenStudy (amtran_bus):

Help me if you can I'm feeling down... Help me get my feet back on the ground!!

OpenStudy (amtran_bus):

Got to fill in all those blanks!

OpenStudy (amtran_bus):

A rectangular storage container with an open top has a volume of 16 m3. The length of its base is twice its width. Material for the base costs $6 per square meter; material for the sides costs $2 per square meter. Express the cost of materials as a function of the width of the base.

OpenStudy (amistre64):

looks like the actual information to fill in is missing from the picture ....

OpenStudy (amtran_bus):

Actually, it should all come from here: A rectangular storage container with an open top has a volume of 16 m3. The length of its base is twice its width. Material for the base costs $6 per square meter; material for the sides costs $2 per square meter. Express the cost of materials as a function of the width of the base.

OpenStudy (amtran_bus):

@amistre64

OpenStudy (yanasidlinskiy):

Wow!!! Have you learned any of this stuff?

OpenStudy (amistre64):



OpenStudy (amistre64):

V = lwh h = V/lw, but l = 2w h = V/2w^2 Ab = l*w As = lh+lh+wh+wh

OpenStudy (amistre64):

well, filling in boxes, if we have an area of lw, and it costs k per ... then the first box has to be what?

OpenStudy (amtran_bus):

The first box, answer is 6.

OpenStudy (amistre64):

good, and they basically give you the next 2 boxes for you

OpenStudy (amtran_bus):

Please excuse me, I am delirious from working this all day. Can you show me those 2? lol sorry.

OpenStudy (amistre64):

2(area of one side) + 2(area of the different side) and there are 2 sides each 2[ 2(area of one side) + 2(area of the different side)]

OpenStudy (amistre64):

how do they define the needed areas?

OpenStudy (amtran_bus):

you mean by h, 2w, w?

OpenStudy (amistre64):

yeah

OpenStudy (amistre64):

2(wh) + 2(2wh) gives us 4 sides .... at $2 each is therefore 2[2(wh) + 2(2wh)] what do we put in the boxes?

OpenStudy (amistre64):

really? 2[2(wh) + 2(2wh)] 2[2(....) + 2(2(.....))] what goes in the boxes?

OpenStudy (amtran_bus):

Duh me.

OpenStudy (amtran_bus):

Yes yes. Sorry about that.

OpenStudy (amistre64):

now the sum of the total cost is the addition of everything we have filled in so far .... so the next 3 boxes are the exact same thing as the first 3 boxes we just did ...

OpenStudy (amtran_bus):

ok. Thanks.

OpenStudy (amistre64):

6(2w^2) + 2[2wh + 2(2wh)] simplify it into : ___w^2 + ___wh

OpenStudy (amistre64):

now they say to eliminate h, use the fact that the volume is equal to 16 m^3, then they ask you to fill in the box that says the volume is equal to _____

OpenStudy (amistre64):

then they say ... divide off the 2w^2, and simplify the results for h

OpenStudy (amtran_bus):

*we are right with everything up to this point* THANK YOU! 16

OpenStudy (amistre64):

then it says to take that h, in terms of w, and fill it into the equation that we got to begin with ...

OpenStudy (amtran_bus):



OpenStudy (amistre64):

have you finished yet? when you do post your stuff and ill post mine

OpenStudy (amtran_bus):

Sorry. This [site] is so slow.

OpenStudy (amtran_bus):

bunch more 12's an 8 or 2, and other stuff. Got it. Thanks.

OpenStudy (amistre64):

:) it was basically alot of filling in what they gave you already :)

OpenStudy (amtran_bus):

U got the deepfreeze in your windows? Thanks again!